OpenStudy (anonymous):

For the interval x > 0 and the function set S = { 3ln(x), ln2, ln(x), ln(5x)} construct a linear ODE of the lowest order

OpenStudy (anonymous):

After taking the wronskian, I get W = 0. Doesn't that mean that there is no linear ODE for this set?

OpenStudy (turingtest):

yeah, it would I believe...

OpenStudy (turingtest):

one of your solutions is ln2 ???

OpenStudy (anonymous):

yeah ln2. I'm not quite certain if I should be excluding that or not, since it can only be differentiated once.

OpenStudy (turingtest):

I don't think that wronskian is zero how did you take the determinant of the 4x4 matrix? method of cofactors?

OpenStudy (turingtest):

oh wait, lnx and 3lnx are linearly dependent, so...

OpenStudy (anonymous):

well the ln2 column gives 3 zeros, right? So I just multiplied -ln2 with the remaining 3x3 matrix The 3x3 matrix is a wronskian of {3/x,1/x,1/(5x)} --- (the derivates of the solutions) Does that sound right so far?

OpenStudy (turingtest):

yeah that is method of cofactors

OpenStudy (anonymous):

let me see if I can get the solution up here

OpenStudy (turingtest):

d/dx(ln(5x))=1/x

OpenStudy (turingtest):

careful with the chain rule there...

OpenStudy (anonymous):

aaahhh that's right! Can't believe I messed that up

OpenStudy (turingtest):

so I get\[W=\left|\begin{matrix}3\ln x&\ln x&\ln(5x)\\\frac3x&\frac1x&\frac1x\\-\frac3{x^2}&-\frac1{x^2}&-\frac1{x^2}\end{matrix}\right|\]looks fishy to me

OpenStudy (turingtest):

rather\[W=-\ln2\left|\begin{matrix}3\ln x&\ln x&\ln(5x)\\\frac3x&\frac1x&\frac1x\\-\frac3{x^2}&-\frac1{x^2}&-\frac1{x^2}\end{matrix}\right|\]

OpenStudy (anonymous):

Won't you have the first row of the new matrix as | 3/x 1/x 1/x | ? Since the original first row with 3lnx etc is multiplied by 0?

OpenStudy (turingtest):

oh shnap, that's why two heads are better than one

OpenStudy (turingtest):

\[W=-\ln2\left|\begin{matrix}\frac3x&\frac1x&\frac1x\\-\frac3{x^2}&-\frac1{x^2}&-\frac1{x^2}\\\frac9{x^3}&\frac3{x^3}&\frac3{x^3}\end{matrix}\right|\]

OpenStudy (anonymous):

Isn't the derivative of -1/x2 = 2/x3 ?

OpenStudy (turingtest):

\[W=-\ln2\left|\begin{matrix}\frac3x&\frac1x&\frac1x\\-\frac3{x^2}&-\frac1{x^2}&-\frac1{x^2}\\\frac6{x^3}&\frac2{x^3}&\frac2{x^3}\end{matrix}\right|\]you know, I really hate arithmetic

OpenStudy (anonymous):

haha who doesn't

OpenStudy (turingtest):

that has to be zero because the last two columns are the same

OpenStudy (anonymous):

Yeah I get the same

OpenStudy (turingtest):

(I'm not actually bothering to do the calculation)

OpenStudy (anonymous):

yeah don't worry about that, the determinant is 0

OpenStudy (anonymous):

It seems like the original set has to be rewritten

OpenStudy (turingtest):

the problem seems to stem from the fact that all the derivatives are linearly dependent constant multiples of each other, (since the constants cancel due to the chain rule) so this thing seems not to want to be a fundamental set of solutions for anything

OpenStudy (turingtest):

it's weird to have one of the solutions be a constant too, I just don't know what to make of that

OpenStudy (anonymous):

Yeah I'm discussing this on another forum at the moment. Someone suggest that I would need to take the smaller set of {ln 2, ln x}